Calculate the molar concentration of calcium, Ca2+ , if 1.2g of calcium phosphate, Ca3(PO4)2 (s), is dissolved in 135mL of solution


1
Expert's answer
2021-01-18T01:56:44-0500

Mr=310 g/mol.

n= m/Mr = 1.2/310 = 0,00387096774 moles.

Cm=n/V = 0,00387096774/0.135 = 0,0286738351 M.